# Over-Squashing in Graph Neural Networks: A Comprehensive survey

### v6

@@ III-A Spatial Graph Rewiring Methods:
Spatial graph rewiring involves modifying the edges of a graph by considering the geometric or spatial relationships between nodes. By optimizing the spatial connections, the rewiring process seeks to improve the overall performance of GNNs in tasks that involve understanding spatial dependencies and relationships among graph elements. The following spatial graph rewiring techniques have been proposed in the literature to mitigate the over-squashing issue, particularly in tasks involving long-range dependencies.
Alon and Yahav [32] were among the first researchers to address the challenge of over-squashing within GNNs. They pointed out that, particularly in prediction scenarios relying on extensive node interactions, GNNs often struggle with over-squashing. This challenge arises because the number of nodes in each node’s receptive field grows exponentially with an increasing number of layers in the network. To address this issue, they proposed a straightforward solution: introducing a fully-adjacent matrix in the final GNN layer. In this method, called the Fully-Adjacent (FA) layer, the GNN layer is reshaped to connect every pair of nodes with an edge. Specifically, in a GNN with $r$ layers, modifications are made only to the $r$-th layer by replacing it with an FA layer. The FA layer ensures that every pair of nodes is directly connected by an edge. Converting an existing layer (1) to be fully-adjacent means that for every node $u\in V$, $N_{u}:=V$ in that layer only. At layer $r$, the transformation can be expressed as:
@@
This transformation does not change the type of layer or add weights but only alters the adjacency of a data sample in a single layer. As a result, the $r-1$ graph layers can still exploit the graph structure using their original sparse topology, while only the $r$-th layer becomes an FA layer, enabling direct interactions among node representations that are aware of the network’s topology and facilitating consideration of nodes beyond their initial neighbors. Topping et al. [35] introduced a novel approach that utilizes graph curvature to address the over-squashing problem in GNNs. Their method, known as Stochastic Discrete Ricci Flow (SDRF), aims to surgically modify negatively-curved edges in order to reduce bottlenecks without significantly altering the statistical properties of the input graph. The core idea of SDRF is to target edges with high negative curvature, which are identified as key contributors to over-squashing. The SDRF algorithm operates in two main steps. First, it identifies edges with minimal Ricci-curvature [^1] Given a graph $G=(V,E)$ with vertices $V$ and edges $E$, the Ricci curvature of an edge $e=(u,v)$ is denoted as $\text{Ric}_{e}$ and is defined in terms of the ratio of the effective resistance between the endpoints of the edge and the length of the edge itself. It can be calculated using various methods, such as discrete Laplacian operators or spectral methods, for details see [51]., indicating potential over-squashing issues. Then, it constructs additional supportive edges around these identified edges to reinforce their structural context within the graph. By strengthening these edges, the algorithm enhances the information flow within the GNN and mitigates the adverse effects of over-squashing.
Black et al. [37] conducted an in-depth analysis of over-squashing by exploring its connection with effective resistance between node pairs, as discussed previously in Section II-B. They proposed a novel graph rewiring method known as the Greedy Total Resistance (GTR) technique to address this issue. The GTR technique aims to minimize the effective resistance between nodes by strategically rewiring the graph. They derived a formula quantifying the impact of adding a specific edge $(u,v)$ on the reduction of total resistance. The total change in resistance when adding an edge $(u,v)$ to a connected graph $G$ with $n$ vertices can be expressed as the difference between the total resistance before and after the addition:
@@
Topping et al. [35] delved into over-squashing, identifying it as the presence of edges with negative curvature. On the other hand, Black et al. and Di Giovanni [37, 34] linked over-squashing to effective resistance or commute time between nodes. In contrast, Gabrielsson et al. [52] proposed a distinct rewiring technique leveraging transformer-inspired positional encoding within a graph framework to expand the receptive field of each node in GNNs. Their method, unlike others, is model-agnostic and focuses on augmenting the original graph with additional nodes and edges using positional encoding as node and/or edge features. This augmentation process does not involve removing existing edges or nodes, ensuring that the original graph remains a subset of the new rewired graph. Specifically, given a graph $G$ and a positive integer $r$, they add edges to all nodes within $r$-hops in $G$ to create a new graph $G_{r}$. Additionally, a fully-connected CLS (classification) node is introduced in $G_{r}$, which connects to all nodes in the graph. However, there’s a concern that in the case of large $r$, the connectivity in the rewired graph $G_{r}$ may become too dense, potentially leading to the loss of the original graph’s topology, a crucial cue for graph-based learning. To address this, they encode the original topology of $G$ into the rewired graph $G_{r}$ using positional encodings [^2] A Positional Encoding refers to a feature that characterizes the spatial arrangement of nodes within a graph, providing information about their global or local positions. These features are often associated with random walk measures and the eigenvectors of the Laplacian matrix [53]., which serve as node and/or edge features. Another approach that probabilistically rewires the graph based on the given prediction task is the Probabilistically Rewired Message-Passing Graph Neural Network (PR-MPNN) [54]. PR-MPNN leverages recent advancements in exact and differentiable k-subset sampling [55] to probabilistically rewire the graph during the learning process. By doing so, PR-MPNN learns to add edges that are relevant to the prediction task while omitting less beneficial ones, effectively addressing issues like over-squashing and under-reaching. Gutteridge et al. [56] propose a novel layer-dependent rewiring technique to address challenges posed by existing rewiring approaches [57, 52] in GNNs. They highlight that some rewiring methods, while aiming to improve connectivity for long-range tasks, compromise the inductive bias provided by graph distance. These approaches enable instant communication between distant nodes at every layer, disrupting the inherent structure. To overcome this, Gutteridge and colleagues introduce a layer-dependent rewiring strategy, dynamically rewired message-passing with delay (DRew) that gradually densifies the graph. This technique allows for enhanced connectivity while preserving the inductive bias provided by graph distance. Additionally, they incorporate a delay mechanism that facilitates skip connections based on both node distance and layer. This mechanism ensures that the graph’s inductive bias is retained, providing a more nuanced and context-aware rewiring approach for improved performance in long-range tasks
@@ III-B Spectral Graph Rewiring Methods:
Deac et al. [47] proposed a model for graph classification task where propagation of information takes place on expander graphs in MPNN and named the method as Expander Graph Propagation (EGP) model. Deac argue that there are many tasks (being graph classification one of them) where only local node level interaction is not sufficient to predict the correct label, specifically for graph classification task where together with local interaction among node feature, global context of graph structure is also important. For such tasks one need to update node features in a manner that is mindful of the global properties of the graph just stacking more message passing layers on the input graph will not sufficient. EGP model strategically employs expander graphs to address challenges associated with bottleneck-induced issues in global information propagation within a graph. Expander graphs are family of graph structures that are fundamentally sparse $(|E|=O|V|)$, while having low diameter: thus, any two nodes in an expander graph may reach each other in a short number of hops, eliminating bottlenecks and oversquashing.
@@
where $\lambda_{1}\geq 0$ represents the first non-zero eigenvalue of the normalized graph Laplacian matrix $\tilde{L}$. This methodology comprises two distinctive layers within a Graph Neural Network (GNN) based in each side of the inequality (9). The first layer, focuses on the left-hand side, the Commute Time Layer (CT-LAYER), serves as a differentiable, parameter-free component tailored to learn the commute time. Meanwhile, the second layer, focuses on the right-hand side, the Gap Layer (GAP-LAYER), functions as the rewiring layer, tasked with optimizing the spectral gap based on the specific network characteristics and task objectives. The CT-LAYER calculates the commute times between nodes and retains edges with significant effective resistance $R_{u,v}$ ($=\frac{\text{Com}(u,v)}{\text{vol}(G)}$) to preserve graph’s topology. Subsequently, the GAP-LAYER adjusts the graph’s adjacency matrix $A$ utilizing ratio-cut and normalized-cut approximations [68] to minimize the spectral gap. During training, both CT-LAYER and GAP-LAYER dynamically learn the weights to predict optimal topology modifications for unseen graphs during testing. This integrated framework endows the GNN with the capability to adaptively learn and implement rewiring strategies, effectively mitigating challenges associated with over-squashing while considering the intricacies of graph structure and task requirements. Banerjee et al. [46] introduced two graph rewiring techniques, namely Random Local Edge Flip (RLEF) and Greedy Random Local Edge Flip (G-RLEF), to alleviate bottlenecks in global information propagation within a graph. RLEF draws inspiration from the flip Markov chain [69, 70, 71], which transforms a connected graph into an expander graph with high probability. For G-RLEF, the authors utilized the relationship between effective resistance $R_{u,v}$ (Eq. 4) and the number of triangles, denoted by $\sharp_{\Delta}(u,v)=|N_{u}\cap N_{v}|$, that contain the edge $(u,v)$:
@@ IV Unifying Approaches to Address Over-Squashing and Over-Smoothing Trade-offs in Graph Neural Networks
Certain methodologies have emerged that tackle the intertwined challenges of over-smoothing and over-squashing in unison, establishing an interconnected relationship between these fundamental limitations within graph neural networks.
@@
where $c_{1},c_{2}$ are positive constant. BORF operates in batches and begins by identifying two sets of edges in each batch: $p$ edges with minimal curvature and $q$ edges with maximal curvature. It then optimizes the graph’s connectivity by adding connections to the minimally curved edges, ensuring efficient communication between distant nodes. This alleviates over-squashing. Additionally, BORF removes the maximally curved edges to prevent over-smoothing, as these can lead to excessive smoothing of node features. Furthermore, the algorithm’s flexibility allows it to operate as a net edge addition, subtraction, or net-zero rewiring, providing adaptability to different data characteristics. In the similar line, Liu et al. [45] connected over-smoothing with edges with positive curvature and over-squashing with negatively curved edges and addressed both problems by focusing on edge removal based on curvature metrics. They harnessed this insight to address the challenges of over-smoothing and over-squashing by introducing a sampling layer driven by Ricci curvature. This sampling layer selectively drops a portion of edges with low Ricci curvature at each GNN layer, effectively mitigating the issues associated with over-smoothing and over-squashing. Recently, a curvature-based edge dropping algorithm known as Curvature-Based Edge Dropping (CBED) is introduced in the work by Dai Shi et al. [75]. This innovative approach strategically removes edges with the highest positive curvature. By doing so, it aims to enhance the model’s adaptability to graphs characterized by heterophily and, in the process, alleviate the issue of over-smoothing.

### v7

@@ I Introduction
Methodology for Literature Review: To ensure a comprehensive and structured review, we followed a systematic methodology to identify relevant literature on over-squashing in Graph Neural Networks (GNNs) and related mitigation techniques. The key steps in our methodology are as follows:
  1. 1.
    Database Selection: We primarily utilized Google Scholar for our search, given its extensive coverage of academic publications across conferences, journals, and preprint servers.
  2. 2.
    Search Strategy: Our search process began by identifying key papers published in top-tier machine learning conferences such as ICML, ICLR, and NeurIPS, which are known for hosting pioneering research on GNNs and related topics.
  3. 3.
    Keyword Selection: To expand the search, we used the following keywords in combination:“over-squashing in GNNs”, “bottleneck effect in GNNs”, “rewiring techniques in GNNs”.
  4. 4.
    Citation Tracking: After identifying foundational papers, we examined their citation networks to discover additional relevant works that extended or improved upon the original methodologies. This backward and forward citation tracking allowed us to include both seminal contributions and recent advancements in the field.
  5. 5.
    Inclusion Criteria: Papers were selected based on their relevance to over-squashing issues in GNNs, proposed solutions such as graph rewiring, and their impact as indicated by citations and peer recognition.
Our systematic approach ensured that the survey encompasses both well-established techniques and emerging methods designed to address over-squashing challenges in GNNs. While there are existing survey articles in the literature that explore specific aspects of this problem, they focus on narrower themes. For instance, the survey by Attali et al. [36] specifically discusses curvature-based rewiring techniques for mitigating over-squashing in GNNs, while Yadati [37] examine over-squashing within the context of hypergraph neural networks. In contrast, our survey provides a more comprehensive overview by consolidating a broader range of methods, including spatial, spectral, and other approaches. To the best of our knowledge, this makes our survey the most extensive resource that offers a holistic perspective on mitigating over-squashing in GNNs.
@@ III-A Spatial Graph Rewiring Methods:
Spatial graph rewiring focuses on modifying graph connections based on geometric or spatial relationships between nodes. By optimizing these spatial connections, the rewiring process aims to improve GNN performance, particularly in tasks involving long-range dependencies. Several spatial graph rewiring techniques have been proposed to mitigate the over-squashing issue.
One of the earliest approaches to address over-squashing in GNNs was introduced by Alon and Yahav [32]. They identified that GNNs struggle with over-squashing in prediction tasks requiring extensive node interactions. As the number of layers increases, the receptive field of each node expands exponentially, making it difficult for information to propagate effectively. To counter this, they proposed a Fully-Adjacent (FA) layer. This method modifies only the final layer in an $r$ layer GNN, replacing it with an FA layer that directly connects every pair of nodes with an edge. Converting an existing layer (1) to be fully-adjacent means that for every node $u\in V$, $N_{u}:=V$ in that layer only. At layer $r$, the transformation can be expressed as:
@@
This transformation retains the original sparse structure in earlier layers while introducing dense connectivity only in the final layer. As a result, the FA layer improves long-range information propagation without altering earlier graph layers. Li et al. [53] introduced a novel rewiring technique called Node-to-Node Distance Relationships (NNDR), designed specifically to mitigate over-squashing. The NNDR method computes higher powers of the adjacency matrix with self-loops, where each entry reflects the number of paths of a given length between node pairs. By analyzing these values, the method identifies nodes prone to over-squashing and strategically introduces new edges to improve information flow. To manage computational complexity, a sampling strategy ensures the rewiring process remains scalable for large graphs. To further enhance the model’s efficiency, the authors introduced a modified GNN architecture called Ordered Neurons for GNNs (O-GNN). Unlike standard relational GNNs (R-GNNs), which aggregate information from all edge types simultaneously, O-GNN distinguishes between original and newly added edges. This separation allows the model to prioritize information from nearby nodes while selectively preserving critical long-range connections. By combining NNDR-based rewiring with O-GNN’s structured learning approach, this method effectively addresses over-squashing while maintaining scalability.
Topping et al. [35] introduced a novel approach that utilizes graph curvature to address the over-squashing problem in GNNs. Their method, known as Stochastic Discrete Ricci Flow (SDRF), aims to surgically modify negatively-curved edges in order to reduce bottlenecks without significantly altering the statistical properties of the input graph. The core idea of SDRF is to target edges with high negative curvature, which are identified as key contributors to over-squashing. The SDRF algorithm operates in two main steps. First, it identifies edges with minimal Ricci-curvature [^1] Given a graph $G=(V,E)$ with vertices $V$ and edges $E$, the Ricci curvature of an edge $e=(u,v)$ is denoted as $\text{Ric}_{e}$ and is defined in terms of the ratio of the effective resistance between the endpoints of the edge and the length of the edge itself. It can be calculated using various methods, such as discrete Laplacian operators or spectral methods, for details see [54]., indicating potential over-squashing issues. Then, it constructs additional supportive edges around these identified edges to reinforce their structural context within the graph. By strengthening these edges, the algorithm enhances the information flow within the GNN and mitigates the adverse effects of over-squashing. Black et al. [39] conducted an in-depth analysis of over-squashing by exploring its connection with effective resistance between node pairs, as discussed previously in Section II-B. They proposed a novel graph rewiring method known as the Greedy Total Resistance (GTR) technique to address this issue. The GTR technique aims to minimize the effective resistance between nodes by strategically rewiring the graph. They derived a formula quantifying the impact of adding a specific edge $(u,v)$ on the reduction of total resistance. The total change in resistance when adding an edge $(u,v)$ to a connected graph $G$ with $n$ vertices can be expressed as the difference between the total resistance before and after the addition:
@@
Topping et al. [35] identified over-squashing as a phenomenon associated with the presence of edges with negative curvature. In contrast, Black et al. and Di Giovanni [39, 34] linked over-squashing to effective resistance or commute time between nodes, offering an alternative perspective on how message-passing bottlenecks arise in graph neural networks (GNNs).
Taking a different approach, Gabrielsson et al. [55] introduced a novel rewiring technique inspired by transformers, utilizing positional encoding to expand the receptive field of each node. Unlike curvature- or resistance-based methods, their model-agnostic approach enhances the original graph by adding nodes and edges while preserving its structure as a subset of the new rewired graph. Given a graph $G$ and an integer $r$, their method constructs a new graph $G_{r}$ by connecting all nodes within $r$-hops and introducing a fully connected CLS (classification) node that links to every node in $G_{r}$. However, a large $r$ can lead to excessive densification, potentially distorting the original topology, which is crucial for graph-based learning. To counteract this, they embed positional encodings into $G_{r}$, retaining information about the structure of $G$ while enabling more effective message passing. These encodings, derived from random walk measures and the eigenvectors of the Laplacian matrix [56], serve as node and/or edge features to maintain structural integrity.
Another model-agnostic strategy for mitigating over-squashing is proposed by Attali et al. [57], who leverage edge curvature to guide information flow dynamically. Their curvature-guided rewiring framework restructures the graph based on edge curvature, enhancing message passing without modifying the underlying message-passing neural network (MPNN) architecture. By introducing a curvature-based homophily metric, the method selectively propagates information along edges with specific curvature properties, distinguishing between positively and negatively curved edges. This flexible approach improves connectivity while reducing bottlenecks. Furthermore, the framework integrates both one-hop and two-hop curvature-based propagation, ensuring efficient long-range communication without unnecessary graph densification. This structured rewiring method not only enhances information diffusion but also remains computationally efficient, making it a promising solution for improving GNN performance.
@@
Another approach that probabilistically rewires the graph based on the given prediction task is the Probabilistically Rewired Message-Passing Graph Neural Network (PR-MPNN) [58]. PR-MPNN leverages recent advancements in exact and differentiable k-subset sampling [59] to probabilistically rewire the graph during the learning process. By doing so, PR-MPNN learns to add edges that are relevant to the prediction task while omitting less beneficial ones, effectively addressing issues like over-squashing and under-reaching. Gutteridge et al. [60] propose a novel layer-dependent rewiring technique to address challenges posed by existing rewiring approaches [61, 55] in GNNs. They highlight that some rewiring methods, while aiming to improve connectivity for long-range tasks, compromise the inductive bias provided by graph distance. These approaches enable instant communication between distant nodes at every layer, disrupting the inherent structure. To overcome this, Gutteridge and colleagues introduce a layer-dependent rewiring strategy, dynamically rewired message-passing with delay (DRew) that gradually densifies the graph. This technique allows for enhanced connectivity while preserving the inductive bias provided by graph distance. Additionally, they incorporate a delay mechanism that facilitates skip connections based on both node distance and layer. This mechanism ensures that the graph’s inductive bias is retained, providing a more nuanced and context-aware rewiring approach for improved performance in long-range tasks.
@@
Spatial Method Node-Node distance based rewiring NNDR [53]
@@
Spatial Method Curvature-based rewiring CCMP [42]
Spatial Method Curvature-based rewiring AFRC [63]
@@
Mixed Method Locality-aware sequential rewiring LASER [65]
@@ III-B Spectral Graph Rewiring Methods:
Deac et al. [49] proposed the Expander Graph Propagation (EGP) model for the graph classification task. In this model, information propagation occurs on expander graphs within an MPNN framework. Deac emphasized that for certain tasks, such as graph classification, relying solely on local node-level interactions is often insufficient for accurate predictions. In these cases, incorporating the global structure of the graph is equally important. Simply stacking additional message-passing layers on the original graph may fail to effectively capture such global information, especially in large graphs where bottlenecks hinder efficient communication. To address this, the EGP model employs expander graphs — a special class of sparse graphs with unique connectivity properties. Expander graphs are characterized by having a number of edges proportional to the number of nodes, i.e., $|E|=O|V|$. This notation emphasizes that while expander graphs maintain sparsity, they achieve remarkably efficient connectivity. Despite having relatively few edges, expander graphs exhibit a low diameter property, meaning any two nodes can be connected in only a few hops. This ensures rapid and efficient information flow across distant nodes, effectively mitigating bottlenecks and reducing the risk of oversquashing. Consequently, the EGP model leverages these properties to improve message propagation in tasks that require both local and global structural awareness.
@@
where $\lambda_{1}\geq 0$ represents the first non-zero eigenvalue of the normalized graph Laplacian matrix $\tilde{L}$. This methodology comprises two distinctive layers within a Graph Neural Network (GNN) based in each side of the inequality (9). The first layer, focuses on the left-hand side, the Commute Time Layer (CT-LAYER), serves as a differentiable, parameter-free component tailored to learn the commute time. Meanwhile, the second layer, focuses on the right-hand side, the Gap Layer (GAP-LAYER), functions as the rewiring layer, tasked with optimizing the spectral gap based on the specific network characteristics and task objectives. The CT-LAYER calculates the commute times between nodes and retains edges with significant effective resistance $R_{u,v}$ ($=\frac{\text{Com}(u,v)}{\text{vol}(G)}$) to preserve graph’s topology. Subsequently, the GAP-LAYER adjusts the graph’s adjacency matrix $A$ utilizing ratio-cut and normalized-cut approximations [74] to minimize the spectral gap. During training, both CT-LAYER and GAP-LAYER dynamically learn the weights to predict optimal topology modifications for unseen graphs during testing. This integrated framework endows the GNN with the capability to adaptively learn and implement rewiring strategies, effectively mitigating challenges associated with over-squashing while considering the intricacies of graph structure and task requirements. Banerjee et al. [48] introduced two graph rewiring techniques, namely Random Local Edge Flip (RLEF) and Greedy Random Local Edge Flip (G-RLEF), to alleviate bottlenecks in global information propagation within a graph. RLEF draws inspiration from the flip Markov chain [75, 76, 77], which transforms a connected graph into an expander graph with high probability. For G-RLEF, the authors utilized the relationship between effective resistance $R_{u,v}$ (Eq. (4)) and the number of triangles, denoted by $\sharp_{\Delta}(u,v)=|N_{u}\cap N_{v}|$, that contain the edge $(u,v)$:
@@
The key distinction between spatial and spectral rewiring lies in how they address bottlenecks in graph structures, particularly in mitigating oversquashing. Spatial rewiring techniques primarily connect nodes within a defined $k$-hop neighborhood [55, 61], or in extreme cases, operate over a fully-connected graph weighted by attention, as seen in Graph-Transformers [78, 56]. While spatial rewiring partially preserves the graph’s locality by restricting connections to nearby nodes or incorporating positional information, these methods often result in denser computational graphs. This increased density raises memory complexity and may inadvertently exacerbate over-smoothing [79, 41]. In contrast, spectral rewiring techniques improve graph connectivity by optimizing graph-theoretic measures related to expansion properties, such as the spectral gap, commute time, or effective resistance [64, 46, 39]. These methods typically introduce fewer edges based on specific optimization criteria, preserving the graph’s sparsity. However, this approach disrupts the locality of the original graph by adding long-range edges between distant nodes. Consequently, while spectral rewiring can enhance connectivity, it may be less effective in addressing oversquashing due to its limited focus on local bottlenecks. To bridge the gap between these two approaches, Barbero et al. [65] proposed the Locality-Aware SEquential Rewiring (LASER) framework. LASER combines the advantages of both spatial and spectral methods by selectively sampling edges to add based on equivariant, optimal conditions. This strategy maintains the inductive bias of spatial rewiring while ensuring the efficiency of spectral techniques, offering a more balanced and edge-efficient solution for mitigating oversquashing.
While spectral methods often rely on global metrics such as the spectral gap, these measures fail to provide insights into the precise locations of local bottlenecks. In contrast, discrete curvature measures offer a localized and computable metric to assess oversquashing directly from the graph’s topology. Building on this idea, Tori et al. [80] explored the effectiveness of curvature-based rewiring in improving performance on real-world graph datasets. Their findings align with Tortorella’s work [81], which evaluated curvature-based rewiring in training-free GNNs and concluded that it seldom provides meaningful benefits for message passing. Tori et al. further demonstrated that curvature-based rewiring often fails to improve performance in real-world datasets. Their analysis attributed previously reported state-of-the-art (SOTA) results to hyperparameter tuning rather than the inherent benefits of the rewiring process itself — a pattern also noted in prior GNN performance studies [82].
@@ IV Unifying Approaches to Address Over-Squashing and Over-Smoothing Trade-offs in Graph Neural Networks
Several methods address the intertwined challenges of over-smoothing and over-squashing in GNNs. Attali et al. [36] provide a detailed review of state-of-the-art rewiring techniques, focusing on their theoretical foundations, practical implementations, and trade-offs, with an emphasis on improving GNN performance in complex settings like heterophilic graphs.
TABLE II: Critical Analysis of Spatial Graph Rewiring Methods
Method Category Pros Cons
Fully-Adjacent (FA) Layer Improves long-range information propagation by introducing dense connectivity in the final layer. Retains the original sparse structure in earlier layers, preserving locality. Dense connectivity in the final layer increases computational complexity. May lead to over-smoothing due to excessive edge additions.
Node-to-Node Distance Relationships (NNDR) Strategically introduces new edges to improve information flow. Scalable due to sampling strategies for large graphs. Computationally expensive due to higher powers of adjacency matrix calculations.
Stochastic Discrete Ricci Flow (SDRF) Targets edges with high negative curvature, reducing bottlenecks. Enhances information flow without significantly altering the graph’s statistical properties. Requires identification of negatively curved edges, which can be computationally intensive.
Greedy Total Resistance (GTR) Minimizes effective resistance between nodes, improving connectivity. Strategically adds edges to maximize reduction in total resistance. Greedy approach may not always find the optimal set of edges.
Positional Encoding Rewiring Expands the receptive field of each node using positional encodings. Preserves the original graph structure while enabling more effective message passing. Large hop counts can lead to excessive densification, distorting the original topology. Positional encodings may not always capture the necessary structural information.
Curvature-Guided Rewiring Dynamically restructures the graph based on edge curvature, enhancing message passing. Reduces bottlenecks by selectively propagating information along specific curvature properties. Computationally expensive due to curvature calculations. Requires careful hyperparameter tuning to balance between positive and negative curvature edges.
Probabilistically Rewired Message-Passing (PR-MPNN) Learns to add task-relevant edges while omitting less beneficial ones. Addresses both over-squashing and under-reaching. Probabilistic rewiring may introduce instability during training. Requires differentiable k-subset sampling, which can be computationally intensive.
Layer-Dependent Rewiring (DRew) Gradually densifies the graph, preserving the inductive bias of graph distance. Skip connections based on node distance and layer improve long-range communication. Dynamic rewiring may increase memory complexity. Requires careful design to balance between connectivity and preserving graph structure.
TABLE III: Critical Analysis of Spectral Graph Rewiring Methods
Method Category Pros Cons
Expander Graph Propagation (EGP) Uses expander graphs to enhance connectivity for long-range tasks. Eliminates bottlenecks by ensuring short paths between nodes. Expander graph construction can be complex and computationally expensive. May disrupt the locality of the original graph by adding long-range edges.
Commute Time and Spectral Gap Rewiring Optimizes the spectral gap to improve graph connectivity. Dynamically learns weights to predict optimal topology modifications. Requires calculation of commute times and spectral gaps, which can be computationally intensive. May not always preserve the original graph’s local structure.
Random Local Edge Flip (RLEF) Transforms the graph into an expander graph with high probability. Improves spectral expansion by flipping edges. Random flipping may not always lead to optimal connectivity. May require multiple iterations to achieve desired spectral properties.
Greedy Random Local Edge Flip (G-RLEF) Expedites spectral expansion by non-uniform sampling of hub edges. Reduces the number of triangles, increasing the spectral gap. Greedy approach may not always find the optimal set of edges to flip. Computationally expensive due to the need to calculate effective resistance and triangle counts.
Locality-Aware Sequential Rewiring (LASER) Combines spatial and spectral rewiring advantages. Selectively samples edges to add based on optimal conditions, preserving locality while enhancing connectivity. Requires careful design to balance between spatial and spectral rewiring. May still introduce some computational overhead due to edge sampling and optimization.
Curvature-Based Rewiring Uses discrete curvature measures to assess over-squashing directly from graph topology. Improves performance by targeting critical structures in the graph. Curvature-based rewiring often fails to improve performance in real-world datasets. Computationally expensive and may require hyperparameter tuning.
First-Order Spectral Rewiring (FoSR) Optimizes the spectral gap by calculating the first-order change caused by adding edges. Reduces over-smoothing by minimizing the Dirichlet energy. Requires calculation of the second eigenvector, which can be computationally intensive. May not always preserve the original graph’s local structure.
Directional Graph Networks (DGN) Propagates information in the direction of eigenvectors, reducing diffusion distance between nodes. Effectively addresses over-squashing and over-smoothing. Requires calculation of eigenvectors, which can be computationally expensive. May not be suitable for all graph structures.
TABLE IV: Critical Analysis of Other Rewiring Methods
Method Category Pros Cons
Maximization-based Graph Convolution (MGC) Uses elementwise maximization to gather information from all powers of the adjacency matrix. Improves information flow across distant nodes, addressing both over-smoothing and over-squashing. May sacrifice some long-range information to alleviate over-smoothing. Requires careful design to balance between aggregation strategies and computational efficiency.
DuoGNN Decouples homophilic and heterophilic interactions, preventing over-smoothing. Captures long-range dependencies while maintaining a compact structure. Requires separate processing of homophilic and heterophilic graphs, increasing complexity. May require careful tuning to balance between different types of interactions.
DeltaGNN Dynamically removes edges with low Interaction Filtering and Clustering (IFC) scores. Restores long-range dependencies by identifying key nodes with high IFC scores. Requires calculation of IFC scores, which can be computationally intensive. May require careful tuning to balance between edge removal and long-range dependency restoration.
Graph Transformers Less susceptible to over-smoothing compared to traditional GNNs. Establishes direct paths between distant nodes, alleviating over-squashing. High computational and memory requirements due to global attention mechanisms. May result in improper training due to blending of local and non-local interactions.
Graph ViT/Mixer MLP Captures long-range dependencies while mitigating over-squashing. Offers improved computational efficiency, speed, and memory advantages. May not be suitable for all graph structures. Requires careful design to balance between local and non-local interactions.
Redundancy-Free GNN (RFGNN) Eliminates redundancy in message propagation using truncated ePaths trees. Balances epath influence, enhancing GNNs’ ability to capture structural information. Requires construction of path-search-trees, which can be computationally intensive. May require careful tuning to balance between epath influence and computational efficiency.
Graph Echo State Network (GESN) Training-free, making it efficient for node classification tasks. Mitigates issues of long-range message passing and over-squashing in heterophilic graphs. Limited to node classification tasks. May not be suitable for all graph structures.
Anti-Symmetric DGN (A-DGN) Preserves long-term dependencies between nodes using anti-symmetric weight matrices. Alleviates over-squashing by enforcing specific properties on the ODE system. Requires careful design to ensure stability and non-dissipative properties. May not be suitable for all graph structures.
Multi-Scaled Heat Kernel GNN (MHKG) Combines smoothing and sharpening effects on node features, managing the trade-off between over-smoothing and over-squashing. Uses curvature-based pooling to guide the pooling process. Requires calculation of heat kernels and curvature values, which can be computationally intensive. May require careful tuning to balance between smoothing and sharpening effects.
Curvature-Based Pooling (CurvPool) Balances over-smoothing and over-squashing by pooling nodes with similar curvature profiles. Retains the original graph structure by remapping old edges to new node clusters. Requires calculation of curvature values, which can be computationally intensive. May require careful tuning to balance between pooling and preserving graph structure.
@@
where $c_{1},c_{2}$ are positive constant. BORF operates in batches and begins by identifying two sets of edges in each batch: $p$ edges with minimal curvature and $q$ edges with maximal curvature. It then optimizes the graph’s connectivity by adding connections to the minimally curved edges, ensuring efficient communication between distant nodes. This alleviates over-squashing. BORF mitigates over-smoothing by removing edges with the highest curvature, as these can excessively smooth node features. The algorithm’s adaptability allows it to function as a net edge addition, subtraction, or net-zero rewiring, making it flexible across different datasets.
In a similar vein, Liu et al. [47] established a connection between over-smoothing and positively curved edges, as well as between over-squashing and negatively curved edges. They tackled both issues by employing curvature-based edge removal, introducing a sampling layer driven by Ricci curvature. This layer selectively removes edges with low Ricci curvature at each GNN layer, effectively mitigating both over-smoothing and over-squashing. However, Fesser and Weber [63] highlight that many curvature-based methods rely on computationally expensive subroutines and require careful hyperparameter tuning, making them impractical for large-scale graphs. To address this, they propose an Augmented Forman-Ricci Curvature (AFRC)-based rewiring technique, which can be computed in linear time. Their method also includes an efficiently computable heuristic for automating hyperparameter selection, eliminating the need for costly tuning. AFRC effectively captures both over-smoothing and over-squashing effects in message-passing GNNs. Dai Shi et al. [86] introduced Curvature-Based Edge Dropping (CBED), a novel algorithm that strategically removes edges with the highest positive curvature. This approach improves the model’s adaptability to heterophilic graphs while simultaneously reducing over-smoothing, further advancing the effectiveness of curvature-based rewiring techniques in GNNs.